OpenStudy (anonymous):

I need someone to help me start this.. I havent done this is some time and I need some help... I know there is an equation for this A car is designed to last an average of 12 years with a standard deviation of 0.8 years. What is the probability that a car will last less than 10 years?

OpenStudy (marigirl):

could this be regarded as a normal distribution question?

OpenStudy (anonymous):

Assume the distribution is normal (Gaussian). Form the z-score, difference between mean and value under consideration, divided by the standard deviation, z = (10-12)/0.8 = -2.5. Look up cumulative normal probability distribution values in a standard table, to find probability of being 2.5 standard deviations less than the mean. Should be somewhat less than 1%.

OpenStudy (anonymous):

Thank you @douglaswinslowcooper but just to make sure I got this right, I am going to ask you a second question but I will give my answer instead of asking for it. Pizza delivery times at Pizza Time are normally distributed with a mean time of 27 minutes and a standard deviation of 3 minutes. Approximately what percent of pizzas are delivered between 24 and 30 minutes? I think it is 32% but I highly doubt it

OpenStudy (anonymous):

Here you have the question of the fraction of a normal distribution that comes within plus or minus a standard deviation of the mean. thus between z=-1 and z= +1, or 0.84-0.16= 0.68 or 68%. You found the fraction delivered outside the 24 to 30 minute time span.